KAGRA MIR (General)
MarcEisenmann - 23:46, Wednesday 05 February 2025 (3902)
restart of pci lab

As we have transformed the previous TAMA storage into a now operational lab space we can now restart the PCI activities.

I've reconnected all required cables, restarted lockin, chopper and laser controllers.

Somehow, the COM port of the translation stage changed from COM6 to COM7.

I've re-set up it on this new COM port (baud rate 9600) with same settings as previously used.

I home the translation stage after moving it by hand outside of the optical components.